OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = -5 and a6 = -5,120?

OpenStudy (anonymous):

first find the common ratio a6 = ar^5 and a1 = -5 so -5120 = 5th root (-5120 / 5) = 5th root 1024 = 4

OpenStudy (anonymous):

the 9th term = ( -5)* 4^8

OpenStudy (anonymous):

D) -524,288

OpenStudy (anonymous):

recall that nth term of a GS is ar^(n-1)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

= -5 * 4 ^ 8

OpenStudy (anonymous):

i found the rght answer it was 327680

OpenStudy (anonymous):

right

OpenStudy (anonymous):

hold on its - 327680 ( * -5)